Work done if 100g water is electrolysed at a constant pressure of 1atm
  1. -19.0 kJ
  2. 203.8 J
  3. 8.33 kJ
  4. 42.20 kJ

Grade:11

1 Answers

Arun
3 years ago
work done = P dV = (m/M) R dT = (50/18) 2 x 25 = 140 calories = 140 x4.2 = 588 J.
 
Hope it helps
 
Thanks and regards